2/3(8x-12)=7-3(x-3)
Please solve the equation.

Answer by mananth(16946)

2/3(8x-12)=7-3(x-3)
.
16x-24 /3 = 7 - 3x+9
16x-24= 3(16-3x)
16x-24=48-9x
25x=72
x=72/25